MIRR A project has an initial cost of $49,675, expected net cash inflows of $9,000 per year for 8 years, and a cost of c...

MIRR

A project has an initial cost of $49,675, expected net cash inflows of $9,000 per year for 8 years, and a cost of capital of 10%. What is the project's MIRR? Round your answer to two decimal places.
Answer #1

Future value = Annuity * [(1 + r)n - 1] / r

Future value = 9000 * [(1 + 0.1)8 - 1] / 0.1

Future value = 9000 * 11.435888

Future value = 102,922.9929

MIRR = (FV / initial investment)1/n - 1

MIRR = (102,922.9929 / 49,675)1/8 - 1

MIRR = 1.0953 - 1

MIRR = 0.0953 or 9.53%